OpenStudy (anonymous):

What is the measure of an acute angle if twice the measure of its supplement is 24 more than five times the measure of its compliment?

OpenStudy (anonymous):

x<90 2(180-x)+24=5(90-x) 360-2x+24=450-5x -2x+24=90-5x 3x=66 x=22

OpenStudy (anonymous):

Thank you so much

OpenStudy (campbell_st):

if the angle is x the supplement is (180 - x) the compliment is (90 - x) the problems reads 2(180 - x) = 5(90 - x) + 24 now solve for x